Solve, by using the method of variation of parameter, the following differential equation
(D^2-3D+2)y = 1/(1+e^-x)
1
Expert's answer
2020-10-25T12:15:01-0400

"\displaystyle

(D^2 - 3D + 2)y = \frac{1}{1 + e^{-x}}\\


\textsf{The auxiliary equation is}\\

m^2 - 3m + 2 = 0\\
(m - 1)(m - 2) = 0\\
m = 1, 2\\


y = y_p + y_c\\


y_c \, \textsf{is the complementary factor}\\

y_p \, \textsf{is the particular Integral}\\

y_c = Ae^x + Be^{2x}\\

y_p = V_1(x)e^{2x} + V_2(x)e^{x}\\


W(x) \, \textsf{is the Wronskian of the solution}\\

W(x) = e^{\int 3 \, \mathrm{d}x}= e^{3x}\\

\begin{aligned}
V_1(x) &= -\int \frac{e^x}{e^{3x}(1 + e^{-x})}\, \mathrm{d}x\\
& = -\int \frac{e^{-2x}}{1 + e^{-x}}\, \mathrm{d}x\\
&= -\int \frac{e^{-x}\cdot e^{-x}}{1 + e^{-x}}\, \mathrm{d}x\\
&= \int \frac{e^{-x}}{1 + e^{-x}}\, \mathrm{d}(e^{-x})\\
&= \int \left(1 - \frac{1}{1 + e^{-x}}\right) \mathrm{d}(e^{-x})\\
&= e^{-x} - \ln(1 + e^{-x}) + C
\end{aligned}\\

\begin{aligned}
V_2(x) &= \int \frac{e^{2x}}{e^{3x}(1 + e^{-x})}\, \mathrm{d}x\\
& = \int \frac{e^{-x}}{1 + e^{-x}}\, \mathrm{d}x\\
&= -\int \frac{1}{1 + e^{-x}}\, \mathrm{d}(e^{-x})\\
&= -\ln(1 + e^{-x}) + C
\end{aligned}\\

\textsf{The constants can be ignored, since we are}\\\textsf{trying to discover non-constant solution to the DE}\\

\begin{aligned}
y &= Ae^{x} + Be^{2x} -e^{x}\ln(1 + e^{-x}) \\&+ e^{2x}(e^{-x} - \ln(1 + e^{-x}))
\end{aligned}\\


y = Ae^{x} + Be^{2x} -e^{x}\ln(1 + e^{-x}) + e^{x} - e^{2x}\ln(1 + e^{-x})\\


\therefore y = Ae^{x} + Be^{2x} - (e^{x}+e^{2x})\ln(1 + e^{-x}) + e^x"